OpenStudy (anonymous):

Get Metal and be Faned The area of a parallelogram is 131.84 square centimeters and the base is 12.8 centimeters. What is the height of this parallelogram?

OpenStudy (whpalmer4):

"medal", "fanned" \[A = b*h\]You know \(A = 131.84\) and \(b = 12.8\), so rearrange the equation to give you \(h\) in terms of \(A, b\) and plug in the numbers.

OpenStudy (anonymous):

131.84-12.8

OpenStudy (whpalmer4):

No, I'm afraid not. If \[A = b*h\], it will not usually be the case that \[h = A-b\]

OpenStudy (whpalmer4):

You need to divide, not subtract. \[A = b*h\]divide both sides by \(b\) because you want \(h\) all by itself: \[\frac{A}{b} = \frac{b*h}{b} = \frac{\cancel{b*}h}{\cancel{b}} = h\] \[h = \frac{131.84}{12.8} =\]

OpenStudy (anonymous):

ok thx so it would be 10.3

OpenStudy (whpalmer4):

Magic 8-ball says "yes"